# Investigating Map-Based Path Loss Models: A Study of Feature Representations in Convolutional Neural Networks

### v1

@@ IV-C Investigating FLIP Performance
On the Canadian data, FLIP demonstrates a turnaround in performance, exhibiting the lowest ensemble RMSE on three of ten holdouts. Similarly to the other two models, FLIP’s RMSE decreased from Table II to Table III, likely for similar reasons. To investigate the causes of FLIP’s successes and failures, we conduct an exhaustive search of features for linear regression where the target is FLIP RMSE per city. We compute the coefficient of determination R2 for all combinations of the following three features: proportion of links having line of sight ($p_{LOS}$), mean total obstruction depth [m] ($\mu_{o}$), and SD of link distance [m] ($\sigma_{d}$). After testing several variables, these three features consistently contributed to high R2 among the various combinations we tried, with diminishing returns when adding more than three. We normalize all features to [0, 1] before performing each regression. We summarize the results in Table IV.
TABLE IV: Features Correlated with FLIP RMSE by City
Features Used in Linear Regression R2 Coefficients
@@
Comparing the signs of the coefficients in Table IV, a higher proportion of line-of-sight links and a higher obstruction depth in a given city both increase RMSE, while having a wider distribution of link distances leads to lower RMSE.
One possible explanation is that the FLIP model achieves low error specifically on non-line-of-sight yet low clutter density links. There are very few line-of-sight links in the UK training set (1.5%), so all three models including FLIP likely struggle with purely line-of-sight links. However, simultaneously, having less clutter may minimize the error imposed by FLIP’s GIS feature extraction described above. Therefore, a small yet non-zero clutter density may minimize FLIP’s error.
@@ V Concluding Remarks
This paper provided a comprehensive analysis of best practices in representing scalar features in deep learning-based path loss models using path profiles. We found that representing simple scalar features using filled image channels can achieve 7.25 dB RMSE in a continentally independent test set of 125 000 measurements. We also examined one model’s strengths and weaknesses after inconsistent RMSE in independent testing.
Future studies will investigate best practices in including supplemental features to further improve the model fit without sacrificing generalization. This may include terrain heights, Fresnel zone information, and clutter types (tree types, building types, etc.). We also plan to further examine the current models’ predictive strengths and weaknesses.

### v2

@@ IV-C Investigating FLIP Performance
On the Canadian data, FLIP demonstrates a turnaround in performance, exhibiting the lowest ensemble RMSE on three of ten holdouts. Similarly to the other two models, FLIP’s RMSE decreased from Table II to Table III, likely for similar reasons. To investigate the causes of FLIP’s successes and failures, we conduct an exhaustive search of features for linear regression where the target is FLIP RMSE per city. We compute the coefficient of determination R2 for all combinations of the following three features: proportion of links having line of sight ($p_{LOS}$), mean total obstruction depth [m] ($\mu_{o}$), and SD of link distance [m] ($\sigma_{d}$). After testing several variables, these three features consistently contributed to high R2 among the various combinations tried, with diminishing returns using more than three. We normalize all features to [0, 1] before each regression. We summarize the results in Table IV.
TABLE IV: Feature(s) Correlated with FLIP RMSE by City
Features Used in Linear Regression R2 Coefficient(s)
$\sigma_{d}$ 0.02 $1.30$
$\mu_{o}$ 0.05 $1.90$
@@
Comparing the signs of the coefficients in Table IV, a higher proportion of line-of-sight links and a higher obstruction depth in a given city both increase RMSE. Conversely, a higher SD of link distance generally decreases RMSE. Though one row contains a positive $\sigma_{d}$ coefficient, the corresponding $R^{2}$ is very low, discrediting the result. One possible explanation for the coefficients is that the FLIP model achieves low error specifically on non-line-of-sight yet low clutter density links. There are very few line-of-sight links in the UK training set (1.5%), so all three models including FLIP likely struggle with purely line-of-sight links. However, the presence of less clutter may reduce the errors imposed by FLIP’s GIS feature extraction described above. Therefore, a small, yet non-zero clutter density may minimize FLIP’s error.
@@ V Concluding Remarks
This paper provided a comprehensive analysis of best practices in representing scalar features in deep learning-based path loss models using path profiles. We found that representing simple scalar features using filled image channels can achieve 7.25 dB RMSE in a continentally independent test set of 125 000 measurements. We also examined one model’s strengths and weaknesses after inconsistent RMSE in independent testing. Future studies will investigate best practices in including supplemental features to further improve the model fit without sacrificing generalization. This may include terrain heights, Fresnel zone information, and clutter types (tree types, building types, etc.). We also plan to further examine the current models’ predictive strengths and weaknesses.